Equipment is purchased at a cost of $39,000. As a result, annual cash revenues will increase by $20,000; annual cash operating expenses will increase by $7,000; straight-line depreciation is used; the asset has a ten-year life; the salvage value is $3,000. Assuming a tax bracket of 34%, determine the accounting rate of return? (round to the nearest %)
Answer #1

a) Account rate of return = Net income/Average investment

Net income = 20000-7000-(39000-3000/10)*66% = 6204

Average investment = (39000+3000/2) = 21000

Accounting rate of return = 6204/21000 = 29.54%

b) Account rate of return = Net income/Initial investment

Net income = 20000-7000-(39000-3000/10)*66% = 6204

Initial investment = 39000

Accounting rate of return = 6204/39000 = 15.91%